# ULSR-GS: Ultra Large-scale Surface Reconstruction Gaussian Splatting with Multi-View Geometric Consistency

### v1

@@ 4.3 Loss Functions
where $\alpha$, $\beta$, are weighting factors that balance the contributions of each error term. $L_{\text{geo}}$ includes two regularization components from 2DGS [11]: depth distortion $\ell_{d}$ and normal consistency $\ell_{n}$. $L_{r}$ represents the RGB reconstruction loss, combining $L_{1}$ loss and the D-SSIM metric as used in 3DGS [12].
Scene LOWER_CAMP CUHK_UPPER LFLS SMBU SZIIT SZTU
Neuralangelo * * * * * * * * * * * * * * * * * * $>$100 h
GOF-60K 0.667 0.652 0.659 0.693 0.672 0.682 0.607 0.659 0.631 0.679 0.784 0.727 0.649 0.695 0.671 - - - 5.3h
PGSR-60K 0.683 0.648 0.665 0.672 0.673 0.672 - - - 0.660 0.795 0.721 0.679 0.642 0.659 0.548 0.542 0.545 5.7h
Ours (4 GPU) 0.681 0.689 0.685 0.690 0.705 0.697 0.684 0.703 0.693 0.737 0.789 0.762 0.705 0.713 0.709 0.713 0.723 0.718 3.7h
Table 1: Quantitative results on GauU-Scene dataset [38]. We report precision, recall, and F-1 score at the threshold $\tau=0.025$. The best, second best, and third best results are highlighted in red, orange, and yellow respectively. ”-” denotes no results due to GPU out of memory. ” *” denotes no result due to no convergence.
Scene Building Rubble Campus Residence Sci-Art
@@
Mega-NeRF 21.48 0.569 0.378 24.70 0.575 0.407 23.93 0.561 0.513 22.86 0.648 0.330 26.25 0.769 0.263
Switch-NeRF 22.07 0.594 0.332 24.93 0.586 0.377 24.03 0.565 0.495 23.41 0.675 0.280 27.07 0.793 0.224
3DGS-60K 22.65 0.742 0.163 26.32 0.786 0.179 23.47 0.703 0.298 23.21 0.791 0.152 24.99 0.805 0.172
2DGS-60K 22.25 0.739 0.198 26.70 0.785 0.204 23.20 0.632 0.342 22.35 0.762 0.162 24.68 0.811 0.219
Mip-Splatting-60K 22.78 0.750 0.160 26.35 0.780 0.172 - - - 23.44 0.785 0.150 24.96 0.810 0.175
Vast Gaussian 23.50 0.804 0.130 26.92 0.823 0.132 26.00 0.816 0.151 24.25 0.852 0.124 26.81 0.885 0.121
Vast Gaussian 23.45 0.780 0.152 26.91 0.805 0.148 25.98 0.793 0.163 24.21 0.822 0.137 26.75 0.878 0.138
ULSR-GS (Ours) 23.49 0.812 0.142 26.95 0.813 0.140 25.99 0.814 0.146 24.24 0.847 0.138 26.81 0.895 0.109
Table 2: Novel view synthesis results on Mill19 [32] and UrbanScene [21] 3D Dataset. The best, second best, and third best results are highlighted in red, orange, and yellow respectively. Vast Gaussian denotes an unofficial reproduction of Vast Gaussian [19]. ”-” denotes no results due to GPU out of memory.
5 Experiments
The comprehensive experimental setups are described in Section 5.1. Assessments of mesh and rendering quality are provided in Section 5.2. Finally, in Section 5.3, we analyze the proposed partition strategy and conduct ablation studies.
5.1 Experimental Setup.
Implementation. In our experiments, we initially employ COLMAP [29] to obtain the SfM results for the entire scene. Subsequently, we align the sparse point cloud using the Manhattan World Alignment, ensuring that the y-axis is perpendicular to the xz plane. We set the angular threshold $\theta_{min}$ = 90, and the distance threshold $D_{max}$= $\sqrt{d_{x}\times d_{y}}$, where $d_{x}$ and $d_{y}$ are the distance of each sub-region. After partitioning, we double the boundary of the initial point cloud for each region for the initialization. This configuration ensures that the edges of each subregion receive sufficient training information, thereby preventing edge mismatches during the stitching process. Redundant points are automatically pruned during training through opacity culling. Experiments for our method are conducted on 4 RTX 4090 GPUs, and each sub-region is trained individually by one GPU.
For each subregion, training is conducted for 40k iterations. Adaptive multi-view densification begin at the 20k-th iteration and terminates at the 30k-th iteration. Regarding the loss function, we set the parameters $\alpha$ = 0.01 and $\beta$ = 0.1. All other loss functions and training parameters are configured identically to those used in 3DGS and 2DGS frameworks. After training, we employ TSDF [35] to extract the mesh of each sub-region then crop using their partition bounding boxes and subsequently stitched together.
@@ 5 Experiments
Datasets and metrics. Our experiments are conducted on large-scale aerial oblique photogrammetry dataset: GauU-Scene (GauU) [38], Mill 19 [32] and UrbanScene3D (US3D)[21]. For evaluating NVS rendering results, we employ PSNR, SSIM, and LPIPS on US3D and Mill 19 datasets. In assessing the reconstructed meshes, we report Precision, Recall, and F-score on GauU-scene dataset. Following the settings of Vast Gaussian [19], we adjust the training configurations for all GS-based methods for a fair comparison.
Compared methods. We compare our proposed ULSR-GS with GS-based methods capable of mesh extraction, including SuGaR [10], 2DGS [11], GOF [47], and PGSR [3]. For NVS rendering evaluation, we include NeRF-based methods Mega-NeRF [32] and Switch-NeRF [25], as well as the GS-based 3DGS [12], Mip-Splatting [46] and Vast Gaussian [19]. Since Vast Gaussian is not open-sourced, we report both the results from the original publication and a GitHub unofficial implementation result.
@@
Figure 5: Qualitative comparison of rendering on the UrbanScene 3D [21] dataset. We demonstrate the zoomed-in surface normal detail of the selected areas in red box. VastGaussian means the unofficial reproduction of Vast Gaussian [19]
5.2 Results
Mesh reconstruction results. Table 1 presents a comparison of reconstruction quality on the GauU-Scene dataset [38]. Our method significantly outperforms other GS-based approaches. When the number of images in the scene (LFLS, SZIIT (see Figure 3), and SZTU) exceeds 1000, the reconstruction quality indicators of 2DGS [11], PGSR [3] and GOF [47] are far behind us. In terms of average training duration, our method exhibits a notably shorter total training time on four GPUs compared to single GPU methods PGSR [3] and GOF [47] at 60K iterations. It is worth mentioning that GOF [47] and PGSR [3] employ similar densification methods, which result in GPU out-of-memory (OOM) issues in certain scenes. This further underscores the advantages of our partitioning strategy. Figure 4 presents several examples of reconstructed meshes, illustrating that our method exceeds existing approaches in both the richness of details and precision. Detailed qualitative and quantitative comparisons are summarised in the supplementary material.
Rendering results. In Table 2, we compare the performance of our proposed ULSR-GS method with various main approaches in different scenes. Although ULSR-GS is designed primarily for precise surface extraction, it also demonstrates exceptional performance on NVS metrics, and Figure 5 presents an representative rendering comparisons.
5.3 Analysis
Point partition. Table 3 presents comparison experiments for partition strategies. Without pair selection (Table 3 (A)), an excessive number of images participate in the training of each region, adversely affecting both the training time and the quality of mesh extraction. These redundant images are often caused by errors in the SfM calculations [29]. To further demonstrate the feasibility of our partitioning strategy, we train ULSR-GS and the baseline method 2DGS [11] on our partition strategies and the Vast Gaussian’s [19] (Table 3 (B), (C) and (D)). The results show that our partition strategy is superior to the image position-based approach in processing large-scale surface extraction tasks. Figure 6 shows a visual comparison on a more challenging close-range dataset with specific path planning [21]. Our method selects a greater number of relevant training images within a smaller area, while the Vast Gaussian [19] strategy selects a wider range, resulting in a lack of information in the partition area. More results can be found in the Supplementary Material.
GauU-Scene (Avg. images: 957) Pre. $\uparrow$ Rec. $\uparrow$ F1 $\uparrow$ Avg. Image Time $\downarrow$
A. w/o sel. (4 GPU) 0.699 0.707 0.702 437 4.5 h
B. Ours + VastG (4 GPU) 0.685 0.700 0.692 359 3.5 h
C. 2DGS + Ours PP (4 GPU) 0.677 0.671 0.674 366 3.0 h
D. 2DGS + VastG PP (4 GPU) 0.663 0.669 0.667 359 2.9 h
E. 2DGS-60K (1 GPU) 0.573 0.631 0.600 957 2.1 h
F. ULSR-GS (4 GPU) 0.713 0.725 0.717 366 3.7 h
Table 3: Partition Strategies Comparison. We conduct experiments on the GauU-Scene dataset [38] scene in 4$\times$4 cells. We reported metrics including Precision, Recall, F-1 Score, Average number of images per cell, and Total training time. VastG PP means the unofficial reproduction partition method of Vast Gaussian [19].
Figure 6: Qualitative comparison of different partition strategy on close-range photogrammetry dataset. We train 2DGS [11] as baseline method on UrbanScene 3D [21] artsci_corase scene.
Ablations. Table 4 showcases ablation experiments for the proposed geometric consistency constraints. Specifically, in Table (4 (A), (B) and (C)) the absence of multi-view depth consistency and normal consistency leads to a decline in mesh quality. We compare densification in Tables 4 (D) and (E), where if we cancel the window mask, the scene will have too many Gaussian points, resulting in OOM. As illustrated in Figure 7, without incorporating $E_{depth}$ and $E_{normal}$, the quality of surface extraction deteriorates, and the edges of each region fail to stitch seamlessly. And removing depth densify results in excessively smooth mesh extraction outcomes.
LOWER_CAMPUS Precision. $\uparrow$ Reccall. $\uparrow$ F-1 Score $\uparrow$
@@ 5.3 Analysis
F. ULSR-GS 0.681 0.689 0.685
Table 4: Geometric Consistency Constraints Ablation Studies. We conducte ablation experiments on the GauU-Scene dataset [38] LOWER_CAMPUS scene in 4$\times$4 cells. We reported metrics including Precision, Recall, F-1 Score. ”-” denotes no data due to GPU out of memory.
@@
6 Conclusion
We present ULSR-GS, a framework dedicated to high- fidelity surface extraction in ultra-large-scale scene. Specifically, our partitioning approach combines with a multi-view selection strategy. Additionally, ULSR-GS employs a multi-view geometric consistency densification to enhance surface details. Experimental results demonstrate that ULSR-GS outperforms other SOTA GS-based works on large-scale benchmark datasets.
ULSR-GS exhibits some limitations when reconstructing highly reflective areas, such as water surfaces and glass buildings. Additionally, despite accurately recalculating rendered depth, computational errors are prone to occur in certain occluded regions. Therefore, future research should explore solutions to address these challenges.
Supplementary Material
7 Detailed Partition Implementations and Comparisons
7.1 Additional implementation details
In this section, we present a description of our partitioning method. As illustrated in Figure 8, we optimize the partition process by focusing on the most relevant parts of the scene.
Figure 8: Density-controlled point cloud filtering for boundary refinement. We demonstrate a case of using density control to filter out sparse point clouds around reconstructed targets from precisely controlled point cloud boundaries in the UrbanScene3D [21] dataset artsci_coarse scene. (a) In the input SfM point cloud, the incorrect matching points outside the reconstructed scene may cause the boundary to become larger. (b) We divide the scene into countless small regions on the xz plane and evaluate the point cloud density of the region, throw out regions that are too low in density. (c) The boundary is significantly reduced by density-controlled SfM point clouds, allowing for fine-grained partition tasks.
Detailed per-point view selection. For each point $P_{k}\in\mathcal{P}$, we consider all possible groups of images that observe $P_{k}$. Specifically, each group $\mathcal{G}_{i}={I_{\text{ref}_{i}},I_{\text{src}_{i}^{1}},I_{\text{src}_{i}^{2}},I_{\text{src}_{i}^{3}}}$ consists of one reference image and three corresponding source images. Our method involves the basic W2C steps. We project $P_{k}$ onto the 2D image planes of each image $I_{j}\in\mathcal{G}_{i}$:
$d_{k_{j}}=\sqrt{(u_{k_{j}}-c_{x_{j}})^{2}+(v_{k_{j}}-c_{y_{j}})^{2}}.$ (18)
Since we want to select the optimal image group for every initial SfM point in the sub-regions. We simply identify the group for $P_{k}$ $\mathcal{G}$ by minimizing $D_{avg}$:
$I_{used}=\bigcup_{P_{k}\in\mathcal{P}}\mathcal{G}_{opt}(P_{k})$ (21)
Sub-regions with insufficient data To address the issue where some sub-regions may contain too few points (especially in boundary areas) after dividing the scene into smaller parts, we evaluate each sub-region as follows: We retain a sub-region only if both the number of points and the number of matched images within it reach at least 10 % of the average per-subregion counts (i.e., the total number of points or images divided by $n^{2}$).
Scene Images Grid Size Used
Campus [21] 5871 $8\times 8$ 41
Residence [21] 2582 $6\times 6$ 20
SZTU [38] 1500 $6\times 6$ 17
LOWER_CAMP [38] 670 $4\times 4$ 12
Table 5: Example Per Scene Partitioning Strategy.
Scene Partitioning Strategy We partition each scene based on the number of images to ensure efficient processing and optimal reconstruction quality. Specifically, we divide the regions according to the following criteria:
  1. 1.
    Scenes with fewer than 1,000 images are divided into a $4\times 4$ grid.
  2. 2.
    Scenes with fewer than 3,000 images are divided into a $6\times 6$ grid.
  3. 3.
    Scenes with more than 3,000 images are divided into an $8\times 8$ grid.
Data Type Five-directional Close-range path planning

### v2

@@ 4.3 Loss Functions
where $\alpha$, $\beta$, are weighting factors that balance the contributions of each error term. $L_{\text{2DGS}}$ includes two regularization components from 2DGS [16]: depth distortion $\ell_{d}$ and normal consistency $\ell_{n}$. $L_{\text{3DGS}}$ represents the RGB reconstruction loss, combining $L_{1}$ loss and the D-SSIM metric as used in 3DGS [19].
5 Experiments
5.1 Implementation
In our experiments, we initially employ COLMAP [44] to obtain the SfM results for the entire scene. Subsequently, we align the sparse point cloud using the Manhattan World Alignment, ensuring that the y-axis is perpendicular to the xz plane. We set the angular threshold $\theta_{min}$ = 90, and the distance threshold $D_{max}$ = $\sqrt{d_{x}\times d_{y}}$, where $d_{x}$ and $d_{y}$ are the distances of each sub-region. After partitioning, we double the boundary of the initial point cloud for each region for the initialization. This configuration ensures that the edges of each subregion receive sufficient training information, thereby preventing edge mismatches during the stitching process. Redundant points are automatically pruned during training through opacity culling. Experiments for our method are conducted on 4 RTX 4090 GPUs, and each sub-region is trained individually by one GPU.
For each subregion, training is conducted for 40k iterations. Adaptive multi-view densification begins at the 20k-th iteration and terminates at the 30k-th iteration. Regarding the loss function, we set the parameters $\alpha$ = 0.01 and $\beta$ = 0.1. All other loss functions and training parameters are configured identically to those used in 3DGS and 2DGS frameworks. After training, we employ TSDF [50] to extract the mesh of each sub-region then crop using their partition bounding boxes and subsequently stitched together.
Figure 3: Examples of experiment datasets. Our experiments are conducted on different types of aerial oblique photography datasets, such as (a) close-range photogrammetry, (b) five-directional (penta-cam) photogrammetry, and (c) hybrid photogrammetry which contains both close-range and penta-cam photogrammetry.
Figure 4: Qualitative comparison with single GPU GS-based methods. We show the full scene comparison with 2DGS [16], GOF[66], and PGSR[5]. From left to right are the LFLS, SZIIT, and SMBU scenes of the GauU-Scene [54] dataset.
Figure 5: Rendering Rensults. We demonstrate the full-scene (zoomed-out) rendering detail of the selected scenes.
5.2 Datasets
As shown in Figure 3, our experiments are conducted on large-scale aerial oblique photogrammetry datasets: GauU-Scene (GauU) [54], UrbanScene3D (US3D) [35], and two custom-collected datasets, namely Scene 1 and Scene 2 (example rendering results can be found at Figure 5). Both Scene 1 (Penta-Cam) and Scene 2 (Close-range + Penta-Cam) contain around 10,000 images.
  1. 1.
@@ 5.2 Datasets
  2. 2.
[…]
  3. 3.
[…]
5.3 Metrics
@@
For evaluating NVS rendering results, we employ PSNR, SSIM, and LPIPS on Urbanscene3D and GauU datasets. In assessing the reconstructed meshes, we report Precision, Recall, and F-score against the GT Lidar pointclouds on GauU-scene dataset, and two custom collected Scene1, and Scene2. Following the settings of Vast Gaussian [29], we adjust the training configurations for all GS-based methods for a fair comparison. The input training photos are downsampled by a factor of 4 (except for the Matrix City dataset [25]).
5.3.1 Rendering quality evaluation
  • PSNR (Peak Signal-to-Noise Ratio): Measures the pixel-wise difference between the rendered images and the ground truth images:
    $\text{PSNR}=10\cdot\log_{10}\left(\frac{\text{MAX}_{I}^{2}}{\text{MSE}}\right)$ (22)
    where is the maximum possible pixel value and MSE is the Mean Squared Error.
  • SSIM (Structural Similarity Index Measure): Evaluates perceived image quality based on luminance, contrast, and structure:
    $\text{SSIM}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$ (23)
  • LPIPS (Learned Perceptual Image Patch Similarity): Measures perceptual similarity by comparing features extracted from pretrained neural networks:
    $\text{LPIPS}(x,y)=\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h,w}\|w_{l}\odot(\hat{y}_{l}^{hw}-\hat{x}_{l}^{hw})\|_{2}^{2}$ (24)
5.3.2 Mesh quality evaluation
Since GauU-scene [54] does not provide a standard evaluation code, we follow the Tanks and Tample [21] evaluation pipeline to quantitatively assess reconstructed meshes against GT LiDAR point clouds:
  1. 1.
  2. 2.
    We uniformly sample number of points from the reconstructed mesh and GT points to 50 million.
  3. 3.
    We utilize KDTreeFlann to compute precision, recall, and F-score using predefined distance thresholds.
  4. 4.
    Precision (Pre): Evaluates the accuracy of reconstructed surfaces against ground truth:
    $\text{Precision}=\frac{|{p\in\mathcal{P}{\text{rec}}|d(p,\mathcal{P}{gt})<\tau}|}{|\mathcal{P}_{\text{rec}}|}$ (25)
  5. 5.
    Recall (Rec): Measures the completeness of reconstruction:
    $\text{Recall}=\frac{|\{p\in\mathcal{P}_{gt}|d(p,\mathcal{P}_{\text{rec}})<\tau\}|}{|\mathcal{P}_{gt}|}$ (26)
  6. 6.
    F-1 Score: Provides a balanced measure combining Precision and Recall:
[…]
5.4 Compared methods.
We compare our proposed ULSR-GS with: (1) single-GPU-based GS methods capable of mesh extraction, including SuGaR [15], 2DGS [16], GOF [66], and PGSR [5]; (2) multi-GPU-based GS method capable of mesh extraction CityGaussianV2 [34]; (3) open source or free-to-use MVS-based photogrammetry projects, including COLMAP and Reality Capture; and (4) we also compare Vast Gaussian’s [29] partiton strategy on rendering tasks and grometry task.
@@ F-1 Score: Provides a balanced measure combining Precision and Recall:
5.5 Experiment results.
5.5.1 Comparison with single-GPU-based GS Methods.
It is important to note that the comparison with single-GPU methods is only as fair as possible. We increase the training iterations of all single-GPU methods by a factor of 2 as much as possible on hardware with 24GB VRAM (a factor of 3 or more will cause most methods to GPU out-of-memory (OOM)), and all training parameters were adjusted accordingly. The full-scene results clearly indicate that our method achieves reconstructions with substantially greater detail and completeness than these baselines.
Figure 4 provides a qualitative comparison of ULSR-GS against three recent single-GPU reconstruction methods: 2DGS [16], GOF [66], and PGSR [5]. In particular, the 2DGS [16] approach produces excessively smooth mesh surfaces that fail to capture fine-grained architectural details, especially in densely built areas. As a result, it cannot accurately reconstruct the intricate geometry of building structures in such regions. The GOF [66] method, which employs a marching tetrahedron algorithm for mesh extraction, often yields meshes with spurious connections to the sky. These artifacts disrupt the continuity of the model and prevent GOF from producing a correct, complete outline of the urban scene. Although PGSR [5] leverages multi-view geometric constraints similar to our method, it selects source views based on those closest in viewpoint to the reference image. In aerial oblique photography scenarios, this strategy frequently chooses source images whose content diverges significantly from the reference view’s observations, hindering effective multi-view geometric consistency enforcement. Consequently, as shown in Figure 7 PGSR’s reconstructions suffer from reduced mesh integrity and completeness. In contrast, our method avoids these issues and preserves both the fine details and the overall structural integrity of the reconstructed urban scene.
Table 2 presents a quantitative evaluation of reconstruction quality on the GauU-Scene dataset [54], using precision, recall, and F1-score metrics at the distance threshold $\tau=0.025$. Our proposed approach demonstrates superior performance across all evaluated scenes, significantly outperforming other single GPU-based Gaussian Splatting (GS) methods in both precision and completeness metrics. Specifically, our method consistently achieves the highest or second-highest scores, reflecting robust reconstruction accuracy and completeness.
Figure 8: Qualitative comparison between ULSR-GS and CityGaussianV2 [34] on the Matrix City dataset [25] Small City scene and the cumtom collected dataset Scene1. Our method is superior to CityGaussianV2 [34] in terms of building integrity and detail preservation.
Regarding computational efficiency, our method, running on four GPUs, achieves shorter training times compared to single-GPU methods GOF [66] and PGSR [5] at 60K iterations. Notably, GOF and PGSR suffer from OOM issues in specific complex scenes due to their similar densification strategies, limiting their applicability for large-scale reconstructions. This clearly highlights the efficacy and scalability of our partitioning strategy, which effectively mitigates GPU memory limitations and enables comprehensive reconstruction even in resource-constrained environments. Furthermore, detailed qualitative visual comparisons in Figure 6 clearly illustrate the superior detail reconstruction capability of our method. The magnified detail results highlight significant improvements in geometric accuracy and fine detail representation compared to the baseline methods.
5.5.2 Comparison with multi-GPU-based GS Methods.
Figure 9: Qualitative comparison between ULSR-GS and Reality Capture and COLMAP on the cumtom collected dataset Scene1. We use a image downsampling factor of 4$\times$ on every method, and use the highest setting of Reality Capture and COLMAP for fair comparison. It can be seen that with the same downsampling factor, ULSR-GS can better reconstruct the thin geometric structure of the scene.
Figure 10: Qualitative evaluation of the Recall metric and the generated DSM.
Figure 11: Qualitative evaluation of different parameter settings. From left to right: GT LiDAR, ULSR-GS with TSDF-fusion at 4x downsampling and 0.4m voxel size, ULSR-GS with TSDF-fusion at 4x downsampling and 0.1m voxel size, CityGaussianV2 with TSDF-fusion at 4x downsampling and 0.4m voxel size, RealityCapture with 4x downsampling and reconstruction quality set to “High Detail”, and RealityCapture with 1x downsampling and reconstruction quality set to “High Detail”.
Table 3 presents a quantitative evaluation comparing our proposed method with existing multi-GPU-based GS methods on the Matrix City dataset [25], and two additional custom-collected datasets (Scene1 and Scene2). Precision, recall, and F1-scores are reported at two different distance thresholds ($\tau=1.0m$ and $\tau=0.5m$).
Our method consistently achieves the best or second-best results across metrics, underscoring the effectiveness and superiority of our per-point partitioning strategy in overcoming scalability bottlenecks, enabling efficient and accurate reconstruction of complex urban scenes. CityGaussianV2 [34], another multi-GPU-based method, similarly requires initial coarse training on a single GPU followed by partitioning for multi-GPU fine-tuning. However, this training paradigm becomes infeasible for datasets comprising nearly 10,000 images, such as Scene1 and Scene2, where the initial sparse point cloud generated by COLMAP exceeds 30 million points, surpassing the memory limitations of a single RTX 4090 GPU with 24GB VRAM. To address this challenge, we applied our proposed per-point partitioning strategy, dividing these two extensive scenes into smaller, manageable 2×2 sub-regions to apply the standard CityGaussianV2 training strategy. Consequently, as shown in the Table 3, this adaptation significantly improves the performance of CityGaussianV2 on the larger scenes.
In the qualitative comparison shown in Figure 8, CityGaussianV2 generates meshes with relatively sharp edges; however, significant holes emerge in smooth and low-texture regions such as roads and ground surfaces, negatively impacting its recall metrics and consequently resulting in lower F1-scores. Additionally, reflective surfaces, such as building glass facades, are inadequately reconstructed by CityGaussianV2, leading to incomplete structural representations. In contrast, our method successfully achieves a balanced reconstruction in both geometric detail and completeness, consistently outperforming CityGaussianV2 in both qualitative visual comparisons and quantitative metrics.
5.5.3 Comparison with MVS-based Methods.
Table 4: Quantitative between MVS-based and GS-based Meshes. We report precision, recall, and F-1 score at distance threshold $\tau=0.5m$. The best, second results are highlighted. "OOM" denotes no results due to VRAM/RAM out of memory during training or TSDF-fusion. "ds" denotes image down sample factor and "v" denotes voxel size. We also report the total time of obtion final mesh results (train + extraction).
Metrics Pre.$\uparrow$ Rec.$\uparrow$ F1$\uparrow$ Face (M) Time (h) $\downarrow$
COLMAP (ds=4) 0.630 0.798 0.704 372.3 29.1
COLMAP (ds=1) 0.707 0.814 0.756 627.9 69.2
Reality Capture (ds=4) 0.629 0.794 0.702 335.6 17.7
Reality Capture (ds=1) 0.703 0.832 0.762 1021.2 61.4
2DGS (ds=4) OOM OOM OOM OOM OOM
CityGSV2 (ds=4, v=0.4m) 0.646 0.753 0.695 357.2 16.1
CityGSV2 (ds=4, v=0.1m) OOM OOM OOM OOM OOM
Ours (ds=4, v=0.4m) 0.632 0.787 0.701 335.6 16.7
Ours (ds=4, v=0.1m) 0.667 0.799 0.727 983.5 19.8
From Table 4, we compare our method against traditional MVS-based approaches (e.g., COLMAP, Reality Capture) and GS-based methods (e.g., 2DGS, CityGSV2). Our approach demonstrates notable advantages in terms of reconstruction efficiency, memory efficiency, and structural detail preservation. As shown in Table 4, our approach achieves a competitive F1 score (0.701 at ds=4, v=0.4m) while maintaining a significantly lower computational cost (16.7 hours) compared to MVS-based methods like Reality Capture, which requires ds=1 to achieve better detail (F1: 0.762, 61.4 hours). Qualitative results in Figure 9 further illustrate that our method preserves fine structures, such as pipes and complex architectural elements, even at ds=4, whereas Reality Capture produces smoother, less detailed meshes. Additionally, recall visualizations in Figure 10 and 11 highlight that our approach generates detailed DSMs at ds=4, whereas MVS-based methods require ds=1 to achieve comparable results, leading to increased computational overhead. Unlike other GS-based methods that suffer from memory limitations at fine voxel resolutions (v=0.1m), our method remains computationally feasible while achieving a high F1 score. These results emphasize the efficiency of our approach in large-scale urban reconstruction, providing high-quality 3D meshes with reduced processing time and memory requirements.
5.6 Ablation Study.
Table 5: Geometric Consistency Constraints Ablation Studies. We conducte ablation experiments on the GauU-Scene dataset [54] in 4$\times$4 cells. We reported metrics including Precision, Recall, F-1 Score. "-" denotes no data due to GPU out of memory.
Precision. $\uparrow$ Reccall. $\uparrow$ F-1 Score $\uparrow$
@@
F. Full 0.681 0.689 0.685
@@
Figure 13: Qualitative comparison on the adapted depth densification ablation studies. Without adaptive densification (e.g. vanilla 2dgs greadient-based densification) results in excessively smooth mesh extraction outcomes and loss rendering quality.
Table 5 presents the results of our ablation experiments, which quantitatively demonstrate the impact of removing geometric consistency constraints on reconstruction quality. Specifically, we observe that the absence of these constraints leads to a notable decline in Precision, Recall, and F1 Score, indicating that geometric consistency is essential for both accuracy and completeness.
5.6.1 Impact of Removing Individual Constraints
When both multi-view depth and normal consistency constraints are removed, it exhibits a noticeable performance decline, reinforcing the importance of geometric constraints in high-quality reconstruction. Interestingly, the F1 score for this condition is slightly higher than that of removing normal consistency alone (Condition C). This suggests that when only normal consistency is absent, the retained depth consistency may introduce conflicting geometric that worsen reconstruction errors, whereas removing both constraints forces the model to rely solely on single image-based information, leading to a different but somewhat more stable degradation. Figure 12 (a) and (b) further highlights the dominant role of multi-view depth and normal consistency in preventing surface misalignment and ensuring smooth reconstructions.
When only the multi-view depth consistency is removed while retaining normal consistency, the moderate decline in performance suggests that although the absence of depth consistency introduces minor geometric misalignment across views, the retained normal consistency still preserves local surface smoothness and alignment, preventing a severe degradation in quality. This result implies that while depth consistency helps refine the global structure, normal consistency alone can still enforce a degree of geometric regularization, maintaining an acceptable level of reconstruction fidelity. When multi-view normal consistency is removed but depth consistency is retained, we observe a substantial drop in performance. This suggests that normal consistency plays a more critical role in maintaining surface quality than depth consistency alone. Without normal consistency, individual surfaces reconstructed from different views may exhibit inconsistent orientations, leading to visible artifacts, non-smooth transitions, and misaligned surface patches.
5.6.2 Impact of Densification Strategies.
When the densification process is removed, Precision and Recall drop slightly to 0.678 and 0.685, respectively, with an F1 Score of 0.682. However, the qualitative comparison in Figure 12 and 13 suggests that densification primarily refines the reconstruction rather than fundamentally altering its structure. The results indicate that while densification improves point distribution and mesh completeness, the geometric consistency constraints (depth and normal consistency) have a more significant impact on the overall reconstruction quality. When remove window mask, ULSR-GS runs out of memory (OOM), preventing evaluation. This suggests that removing the window mask leads to an excessive increase in Gaussian points, which significantly inflates the computational burden. The failure to complete the reconstruction in this setting highlights the importance of an adaptive sparsification mechanism, which prevents unnecessary computational overhead while preserving reconstruction quality.
5.6.3 Impact of Partition Strategies
(a) Geometry Reconstruction
Table 6 presents comprehensive comparative experiments analyzing different partitioning strategies. In the absence of optimized pair selection (as demonstrated in Table 6(A)), a significant number of redundant images are included in training for each region. This redundancy not only substantially increases computational overhead but also negatively impacts the quality and precision of the extracted meshes. These redundant images typically originate from inaccuracies inherent to SfM computations [44], which fail to effectively filter images based solely on geometric consistency.
To further illustrate the robustness and effectiveness of our proposed partitioning strategy, we conducted experiments comparing our strategy against the image location-based partitioning method employed by Vast Gaussian [29]. Specifically, we evaluated both ULSR-GS and the baseline method 2DGS [16] under different partitioning conditions (Table 6(B), (C), and (D)). Our results demonstrate a clear superiority of our method, particularly in large-scale surface reconstruction scenarios.
Notably, Figure 15 visually illustrates the effectiveness of our partitioning strategy using a more challenging close-range oblique photogrammetry dataset featuring intricate and specific UAV flight paths [35]. In this dataset, our approach successfully identifies and selects images highly relevant to the reconstruction region, significantly enhancing local detail preservation and accuracy. Conversely, the partitioning strategy of Vast Gaussian, which relies strictly on the projection of camera positions onto a 2D plane, selects images from a broader area without adequate precision. Consequently, critical images positioned outside of a predefined boundary, yet crucial for reconstructing complex structures, are erroneously excluded. Simultaneously, the inclusion of unrelated, distant images introduces noise, ultimately deteriorating the reconstruction quality. Additionally, on datasets collected using multi-directional Penta Cam setups in Figure 14, where camera paths and orientations are highly irregular, our point-to-photo partitioning approach consistently outperforms the camera projection-based method.
Figure 14: Qualitative comparison of partition strategies on a penta-cam dataset. The black bounding box is the same partition area. The image partition boundary for VastGaussian and and the point cloud partition boundary for ULSR-GS ars set to the same size according to the black boundary. Below are the meshes after clipping to the box size.
Figure 15: Qualitative comparison of partition strategies on a close-range dataset. The black bounding box is the same partition area. The image partition boundary for VastGaussian and and the point cloud partition boundary for ULSR-GS ars set to the same size according to the black boundary. Below are the meshes after clipping to the box size.
(b) Novel view synthesis.
Table 7 quantitatively compares different partitioning methods, specifically our point-based partitioning approach and the image-based method employed by Vast Gaussian [29]. The results indicate that the Vast Gaussian partitioning strategy is more suitable for rendering tasks. This is because our method prioritizes the reconstruction of the scene’s primary structure by excluding the distant sparse SfM points during sub-region partitioning. Figure 16 provides a qualitative comparison to further illustrate this difference. When training ULSR-GS using different partitioning strategies, our partition method shows a trade-off: while it loses performance in reconstructing the background, it preserves significantly more detail in the reconstruction of the primary scene components.
Table 6: Partition Strategies Comparison. We conduct experiments on the GauU-Scene dataset [54] scene in 4$\times$4 cells. We reported metrics including Precision, Recall, and F-1 Score, Average number of images per cell, and Total training time. VastG means the unofficial reproduction partition method of Vast Gaussian [29].
GauU-Scene (Avg. images: 957) Pre. $\uparrow$ Rec. $\uparrow$ F1 $\uparrow$ Avg. I. Time $\downarrow$
A. Ours w/o View Selection 0.699 0.707 0.702 437 4.5 h
B. Ours + Vast-G 0.685 0.700 0.692 359 3.5 h
C. 2DGS + Ours 0.677 0.671 0.674 366 3.0 h
D. 2DGS + Vast-G 0.663 0.669 0.667 359 2.9 h
E. 2DGS-60K (1 GPU) 0.573 0.631 0.600 957 2.1 h
F. ULSR-GS (4 GPU) 0.713 0.725 0.717 366 3.7 h
Data Type Five-directional Close-range